Can you please help on this question. The length of a rectangular fish tank is twice its witdth. a second tank, which is 5 inches longer and 2 inches narrower than the first tank, has a perimeter of 120 inches. Find the length and width of each fish tank.
i need help on this promble which i do not get and need help.

Answer by lwsshak3(11628)

The length of a rectangular fish tank is twice its width. a second tank, which is 5 inches longer and 2 inches narrower than the first tank, has a perimeter of 120 inches. Find the length and width of each fish tank.
***
let x=width of 1st fish tank
2x=length of 1st fish tank
..
x-2=width of 2nd fish tank
2x+5=length of 2nd fish tank
2(width+length)=perimeter
2(x-2+2x+5)=120
6x+6=120
6x=114
x=19
2x=38
x-2=17
2x+5=43
width of 1st fish tank=19 inches
length of 1st fish tank=38 inches
width of 2nd fish tank=17 inches
length of 2nd fish tank=43 inches